OpenStudy (ahsome):

Find a and b if (a+bi)^2=12+16i. I don't know what to do?

OpenStudy (ahsome):

I have expanded and got: \[ab=8\] and \[a^2-b^2=12\]

hartnn (hartnn):

ab =8 so a=8/b plug this in a^2-b^2 =12 and you will get an equation only in a

OpenStudy (ahsome):

I need to find the real and imaginary value

hartnn (hartnn):

(8/b)^2 - b^2 =12 64 - b^4 = 12b^2 take b^2 = x x^2 +12x-64 =0 solve this quadratic to get b^2

OpenStudy (ahsome):

Then I can square root that answer to get b, put that in the ab=8 to get the a value, right?

hartnn (hartnn):

there's an alternative way too, if you know how to convert complex no. from cartesian to polar co-ordinates

OpenStudy (ahsome):

Nope, don't know that way

hartnn (hartnn):

ok, so go with "Then I can square root that answer to get b, put that in the ab=8 to get the a value, right?" yes thats right

OpenStudy (ahsome):

Great. But what is the real and imaginary number? I only get A and B that way, not real and imaginary.

hartnn (hartnn):

"Find a and b " you want a and b only

OpenStudy (ahsome):

Sorry, wrote the question wrong. I want to find the real and imaginary number. I guessed that was the A and B value, right?

hartnn (hartnn):

'the real and imaginary number' of what ? Post the exact question you were asked...

hartnn (hartnn):

'(a+bi)^2=12+16i , find the real and imaginary number' this does not make sense

OpenStudy (ahsome):

K: Find the real numbers \[a\] and \[b\] if \[(a+bi)^2=12+16i\]

hartnn (hartnn):

YES, a and b are both real numbers

hartnn (hartnn):

"bi" is an imaginary number

hartnn (hartnn):

and 'a+bi' is a complex number

OpenStudy (ahsome):

OH, Now I get it. Thanks (sorry for being bothersome)

hartnn (hartnn):

no problem :) welcome ^_^